# The Impact of Partial Computations on the Red-Blue Pebble Game

### v1

@@ Abstract.
We study an extension of the well-known red-blue pebble game (RBP) with partial computation steps, inspired by the recent work of Sobczyk (Sobczyk, 2024). While the original RBP assumes that we need to have all the inputs of an operation in fast memory at the same time, in many concrete computations, the inputs can be aggregated one by one into the final output value. These partial computation steps can enable pebbling strategies with much smaller I/O cost, and in settings where such a step-by-step aggregation is possible, this extended red-blue pebble game offers a much more realistic cost model.
@@
ccs: Theory of computation Models of computationccs: Computer systems organization Processors and memory architectures©Pál András Papp, Aleksandros Sobczyk and Albert-Jan N. Yzelman, 2025. This is the author’s full version of the work, posted here for personal use. Not for redistribution. The definitive version was published in the 37th ACM Symposium on Parallelism in Algorithms and Architectures (SPAA 2025), https://doi.org/10.1145/3694906.3743320.
@@ 1. Introduction
The optimal execution of a complex computational task with limited resources is a fundamental question in computer science, with applications ranging from scientific simulations to machine learning. In many of these applications, the main performance bottleneck is in fact the vertical data movements between different layers of memory, known as I/O operations. One of the prominent tools to analyze this problem is the red-blue pebble game of Hong and Kung (Hong and Kung, 1981), which captures the I/O complexity of a computation in a two-level memory hierarchy: it studies the amount of vertical communication required when we want to execute the computation using a fast memory of limited size (e.g. cache), and a slow memory of much larger capacity (e.g. RAM). This model has been a vital tool to measure I/O complexity in different computations, and has been studied extensively in the past decades.
@@
A pebbling consists of a sequence of rule applications such that finally, all the sink nodes of $G$ (i.e. the outputs of the computation) have a blue pebble. Furthermore, we require that at any point during the pebbling, the number of red pebbles placed on $G$ is at most $r$, i.e. the fast memory capacity is not exceeded. The cost of a pebbling strategy is defined as the total number of save and load operations executed during the pebbling; intuitively, the compute and delete steps are considered free; indeed, these are often significantly cheaper in practice.
@@
One fundamental property of RBP is that it requires all inputs of an operation to be collected into fast memory at the same time. Specifically, if a node $v$ has $d$ incoming edges in our DAG, then any valid pebbling in RBP requires at least $(d\!+\!1)$ red pebbles simultaneously to compute this node: $d$ pebbles on the in-neighbors of $v$, and one more on $v$. Hence a valid pebbling only exists if $r\!\geq\!(\Delta_{in}\!+\!1)$.
@@
In a plethora of practical cases, however, the subtasks in our computation are associative and commutative operators, e.g. the addition of scalars or matrices, applied on a larger set of inputs. In this case, the separate inputs could be added to the aggregated value in any specific order, or if desired, we could even decide to first aggregate e.g. only half of the inputs of a node $v$, store the temporary result in (fast or slow) memory, then do some computations in another part of the DAG, and finally continue processing the remaining inputs of $v$.
@@
The idea of incorporating such partial computation steps into RBP was recently raised by Sobczyk (Sobczyk, 2024). Allowing partial computations not only enables us to analyze I/O costs when $r<(\Delta_{in}\!+\!1)$, but it can also drastically reduce the number of required I/O steps for DAGs with $r\geq(\Delta_{in}\!+\!1)$. Thus, while this extension restricts the scope of RBP to computational DAGs with specific kind of operators, it provides a much more accurate model of the actual I/O cost of the computation in these domains.
@@
In this paper, we provide a comprehensive analysis of the impact of such partial computations on the red-blue pebble game. We begin by discussing the fundamental properties of this partial-computing red-blue pebble game (PRBP), and showing some example DAGs where allowing partial compute steps indeed decreases the number of required I/O operations. We then analyze how the optimum cost on a given DAG differs in RBP and PRBP in general. In particular, we show that partial computations can decrease the required I/O cost by up to a linear factor, but in a general DAG, it is NP-hard to decide whether they reduce the optimum cost at all.
@@
We then focus on developing lower bounds on the I/O cost of specific computations in PRBP. We first show in a counterexample that lower bounds derived from $S$-partitioning, which is a standard tool for RBP, do not directly carry over to PRBP. Instead, we discuss two different ways to modify the $S$-partitioning concept such that the resulting bounds also apply to PRBP. We then use these newly developed tools to derive I/O lower bounds in PRBP for some paramount computational tasks, such as fast Fourier transform, matrix multiplication, and self-attention in neural networks. These are all extensively studied computations with known optimal I/O complexity, and hence, unsurprisingly, the obtained PRBP lower bounds are identical to those in RBP, i.e. partial computations do not allow for a lower I/O cost in these specific DAGs. Nonetheless, the formal proof of this fact is still a crucial step towards understanding the properties of these problems, and I/O complexity in general. Furthermore, the potential impact of our new tools extends beyond these classic, well-structured computations: they may also be used in the future to study lower bounds for irregular graphs or sparse computations.
@@ 2. Related Work
In general, pebble games on graphs have been used before to capture various different aspects of computation, such as time-memory trade-offs, non-determinism or reversibility (Cook, 1973; Bennett, 1989; Cook and Sethi, 1974).
@@
The red-blue pebble game was introduced by Hong and Kung to analyze the I/O complexity of a computational task (Hong and Kung, 1981). Their seminal paper also present $S$-partitionings, a vital tool to derive corresponding I/O lower bounds on specific structured DAGs. $S$-partitions were used to establish lower bounds for various computations both in (Hong and Kung, 1981) and following works (Elango et al., 2015; Ranjan et al., 2012). Other papers have used different methods to derive I/O lower bounds on concrete computations (Jain and Zaharia, 2020; Gleinig and Hoefler, 2022; Scott et al., 2015). Recently, there is also a growing interest in using RBP to derive bounds on the I/O cost of crucial operations in neural networks (Saha and Ye, 2024; Li et al., 2024).
@@
Besides lower bounds, another natural direction of work is the computational complexity of RBP. This includes NP-hardness of different model variants (Demaine and Liu, 2018; Papp and Wattenhofer, 2020; Blocki et al., 2018) and inapproximability results (Böhnlein et al., 2025), as well as approximation algorithms for special cases (Gleinig and Hoefler, 2022; Carpenter et al., 2016).
@@
There are also several papers that discuss the idea of generalizing the RBP model to multiple processors or to deeper memory hierarchies (Böhnlein et al., 2025; Carpenter et al., 2016; Savage, 1995; Savage and Zubair, 2008; Elango et al., 2014). However, these settings are typically very complex, and thus the theoretical results are limited to concrete DAGs or special cases.
@@
A recent preprint (Sobczyk, 2024) raises the idea of extending RBP with partial computations. Our work is inspired by this approach, but rather different both in terms of the model and the direction of the results. In particular, the model of (Sobczyk, 2024) assumes that all nodes have an initialization value that always needs to be loaded from slow memory, motivated by e.g. evaluating polynomials. This incurs an extra I/O step for each sink non-sink node, and hence results in a significantly different (often much higher) I/O cost for most DAGs than either the original RBP or our model. Due to this, most claims in our paper do not carry over to the setting of (Sobczyk, 2024). In contrast to this, our PRBP model is a direct extension of RBP, where any RBP pebbling strategy translates to a partial-computing PRBP solution of the same I/O cost.
@@
In terms of results, (Sobczyk, 2024) mainly focuses on DAGs that are not even computable in RBP due to $r\!<\!(\Delta_{in}\!+\!1)$, and shows NP-hardness and approximation methods for the edge case when the cache size is only $r\!=\!2$. In contrast, we focus on understanding the properties of PRBP in general, and its relation to RBP, for example, how the optimal I/O costs differ between the two models, and which general results from RBP carry over to PRBP.
@@
Regarding the actual model definition, we also change the terminology from (Sobczyk, 2024) slightly to align it with the original RBP terminology. Specifically, we use “light red” and “dark red” pebbles to denote the two different states of values kept in cache, to signify that these all correspond to red pebbles in the original RBP. Furthermore, the incoming edges of a node that are already aggregated into the final value are not deleted, but rather marked in our case; this indicates that the structure of the computation remains unchanged, and also allows to extend the setting to re-computations.
@@ 3. Pebbling With Partial Computations
In the partial-computing red-blue pebble game (PRBP), we have two different kinds of red pebbles: a light red pebble on node $v$ indicates that the current value of $v$ is also up-to-date in slow memory, whereas a dark red pebble indicates that the newest value of $v$ is only present in fast memory, and thus, if ever needed in the future, it has to be first saved to slow memory before deletion. The total number of light and dark red pebbles on the DAG is again bounded at any point by the fast memory size $r$. Blue pebbles still represent values in slow memory, and their number is unlimited.
@@
    partial compute: consider an unmarked edge $(u,v)$ such that: (i) all incoming edges of $u$ are marked, (ii) there is a (light or dark) red pebble on $u$, and (iii) there is either a (light or dark) red pebble on $v$, or no pebble on $v$ at all. Then replace all pebble(s) on $v$ by a dark red pebble, and mark the edge $(u,v)$.
@@
Similarly to RBP, the initial state only has blue pebbles on the source nodes. However, in the terminal state, we now not only require that every sink has a blue pebble, but also that every edge of the DAG is marked. At any point, the number of red pebbles must be at most $r$. As in RBP, the cost of a pebbling is defined as the total number of save and load operations.
Note that in contrast to RBP, this game allows a valid pebbling for any DAG with as few as $r\!=\!2$ red pebbles, by considering a topological order of $G$ and marking each incoming edge of the next node, with the inputs loaded and outputs saved when necessary.
In our paper, we consider the one-shot version of RBP and PRBP, which means that during a pebbling, the compute rule can be applied only once for every node $v\!\in\!V$ in RBP, and only once for every edge $(u,v)\!\in\!E$ in PRBP. One-shot RBP is a prominent version in the literature (Papp and Wattenhofer, 2020; Carpenter et al., 2016; Elango et al., 2014, 2015); we focus on this variant because allowing so-called re-computation steps in PRBP would raise further modeling questions. In Section 8.1, we briefly discuss other variants of RBP, and how the PRBP model extends to these settings.
@@ 4. Fundamental properties of PRBP
One fundamental question about this extended model is how the partial compute steps affect the optimal I/O cost of a specific computation. For a given DAG $G$ and choice of $r$, let us use $\texttt{OPT}_{RBP}$ and $\texttt{OPT}_{PRBP}$ to denote the cost of the optimal pebbling strategy in RBP and PRBP, respectively. We will only compare these metrics for concrete cases when $(\Delta_{in}+1)\leq r$, i.e. a valid pebbling exists in both RBP and PRBP. Regarding notation, we will also use $[d]$ to denote the set of integers $\{1,...,d\}$.
@@
Furthermore, note that both in RBP and PRBP, any valid pebbling strategy needs to load every source node at least once, and needs to save every sink node at least once. If the total number of source and sink nodes in $G$ is $m$, then we refer to this value $m$ as the trivial cost, and we have $\texttt{OPT}_{RBP}\geq m$ and $\texttt{OPT}_{PRBP}\geq m$.
@@ 4.1. Motivating examples
Towards a more practical example, consider a a fundamental operation in linear algebra, namely, the multiplication $A\!\cdot\!x=y$ of an $m\times m$ matrix $A$ with an $m\times 1$ vector $x$. The corresponding computational DAG consists of $m^{2}\!+m$ source nodes, $m^{2}$ intermediate nodes with in-degree $2$, and $m$ sink nodes of in-degree $m$.
@@
In PRBP, we can devise a strategy that keeps the $m$ partially computed values of the output $y$ in fast memory, and only uses $3$ further red pebbles besides this. In particular, we can iterate through the entries $x_{i}$ of the input vector, first loading $x_{i}$ into fast memory, and then for each row of the matrix, we load the corresponding $A_{j,i}$, multiply it with $x_{i}$, and add this $y_{j}$. Note that $A_{j,i}$ can be deleted after computing $A_{j,i}\!\cdot\!x_{i}$, and then $A_{j,i}\!\cdot\!x_{i}$ can be deleted after updating $y_{j}$, so this only requires two red pebbles besides the one on $x_{i}$. Afterwards, $x_{i}$ can also be deleted. This strategy comes with only the trivial I/O cost of $\texttt{OPT}_{PRBP}=m^{2}+2m$.
In RBP, we can show that at least one non-trivial I/O step always happens between the computation of any two consecutively computed sink nodes $y_{j_{1}}$ and $y_{j_{2}}$. Note that computing $y_{j_{1}}$ requires $(m+1)$ red pebbles simultaneously, and thus we can have at most $(m-1)$ red pebbles in the remaining nodes. As such, there is an index $i\in[m]$ such that neither the source node $x_{i}$ nor the intermediate node $A_{i,j_{2}}$ has a red pebble. If $A_{i,j_{2}}$ already has a blue pebble, then this means we have to load it before computing $y_{j_{2}}$. Otherwise, if $A_{i,j_{2}}$ has no pebble, then we need to load $x_{i}$ to compute it; note that $x_{i}$ has been loaded before to compute $A_{i,j_{1}}$, so this is also a non-trivial I/O. Altogether, this gives $\texttt{OPT}_{RBP}\geq m^{2}\!+\!3m-1$. ∎

### v2

@@ Abstract.
We study an extension of the well-known red-blue pebble game (RBP) with partial computation steps, inspired by the recent work of Sobczyk (partial). While the original RBP assumes that we need to have all the inputs of an operation in fast memory at the same time, in many concrete computations, the inputs can be aggregated one by one into the final output value. These partial computation steps can enable pebbling strategies with much smaller I/O cost, and in settings where such a step-by-step aggregation is possible, this extended red-blue pebble game offers a much more realistic cost model.
@@
$c$s: Theory of computation Models of computationccs: Computer systems organization Processors and memory architectures©Pál András Papp, Aleksandros Sobczyk and Albert-Jan N. Yzelman, 2025. This is the author’s full version of the work, posted here for personal use. Not for redistribution. The definitive version was published in the 37th ACM Symposium on Parallelism in Algorithms and Architectures (SPAA 2025), https://doi.org/10.1145/3694906.3743320.
@@ 1. Introduction
The optimal execution of a complex computational task with limited resources is a fundamental question in computer science, with applications ranging from scientific simulations to machine learning. In many of these applications, the main performance bottleneck is in fact the vertical data movements between different layers of memory, known as I/O operations. One of the prominent tools to analyze this problem is the red-blue pebble game of Hong and Kung (RBpebbling1), which captures the I/O complexity of a computation in a two-level memory hierarchy: it studies the amount of vertical communication required when we want to execute the computation using a fast memory of limited size (e.g. cache), and a slow memory of much larger capacity (e.g. RAM). This model has been a vital tool to measure I/O complexity in different computations, and has been studied extensively in the past decades.
@@
A pebbling consists of a sequence of rule applications such that finally, all the sink nodes of $G$ (i.e. the outputs of the computation) have a blue pebble. Furthermore, we require that at any point during the pebbling, the number of red pebbles placed on $G$ is at most $r$, i.e. the fast memory capacity is not exceeded. The cost of a pebbling strategy is defined as the total number of save and load operations executed during the pebbling. Intuitively, the compute and delete steps are considered free; indeed, these are often significantly cheaper in practice.
@@
One fundamental property of RBP is that it requires all inputs of an operation to be collected into fast memory at the same time. Specifically, if a node $v$ has $d$ incoming edges in our DAG, then any valid pebbling in RBP requires at least (d+1)(d\!+\!1) red pebbles simultaneously to compute this node: $d$ pebbles on the in-neighbors of $v$, and one more on $v$. Hence a valid pebbling only exists if $r\!\geq\!(\Delta_{in}\!+\!1)$.
@@
In contrast to this, in many practical cases, the subtasks in our computation are associative and commutative operators, e.g. the addition of scalars or matrices, applied on a larger set of inputs. In this case, the separate inputs could be added to the aggregated value in any specific order, or if desired, we could even decide to first aggregate e.g. only half of the inputs of a node $v$, store the temporary result in (fast or slow) memory, then do some computations in another part of the DAG, and finally continue processing the remaining inputs of $v$.
@@
The idea of incorporating such partial computation steps into RBP was recently raised by Sobczyk (partial). Allowing partial computations not only enables us to analyze I/O costs when $r<(\Delta_{in}\!+\!1)$, but it can also drastically reduce the number of required I/O steps for DAGs with $r\geq(\Delta_{in}\!+\!1)$. Thus, while this extension restricts the scope of RBP to computational DAGs with specific kind of operators, it provides a much more accurate model of the actual I/O cost of the computation in these domains.
@@
In this paper, we extend the red-blue pebble game with such partial computations. We begin by discussing the fundamental properties of this partial-computing red-blue pebble game (PRBP), and showing some example DAGs where allowing partial compute steps indeed decreases the number of required I/O operations. We then analyze how the optimum cost on a given DAG differs in RBP and PRBP in general. In particular, we show that partial computations can decrease the required I/O cost by up to a linear factor, but in a general DAG, it is NP-hard to decide whether they reduce the optimum cost at all.
@@
We then focus on developing lower bounds on the I/O cost of specific computations in PRBP. We first show in a counterexample that lower bounds derived from $S$-partitions, which is a standard tool for RBP, do not directly carry over to PRBP. Instead, we discuss two different ways to modify the $S$-partition concept so that the resulting bounds also apply to PRBP. We then use these newly developed tools to derive I/O lower bounds on some crucial computational tasks, such as fast Fourier transform, matrix multiplication, or self-attention in neural networks. Since the optimal solutions to these problems have been studied extensively, and their I/O complexity matches known lower bounds from RBP, it is not surprising that the obtained PRBP lower bounds are identical to those in RBP, i.e. partial computations do not improve upon the state-of-the-art in these DAGs. Nonetheless, our formal proof of this fact is still a crucial step towards understanding the properties of these problems, and I/O complexity in general.
@@ 2. Related Work
The concept of pebble games on graphs has been widely used to capture various different aspects of computing, such as time-memory trade-offs, non-determinism or reversibility (cook1973observation; otherpebbling2; otherpebbling3).
@@
The red-blue pebble game was introduced by Hong and Kung to analyze the I/O complexity of a computational task (RBpebbling1). Their seminal paper also presents $S$-partitions, a vital tool to derive corresponding I/O lower bounds on specific structured DAGs. These $S$-partitions were used to establish lower bounds for various computations both in (RBpebbling1) and in several following works (elango2015characterizing; ranjan2012upper). Other papers have used different methods to derive I/O lower bounds on concrete computations (jain2020spectral; RBpebbling5; scott2015matrix). Recently, there is also a growing interest in using RBP to derive bounds on the I/O cost of crucial operations in modern neural networks (ml1; ml2).
@@
Besides lower bounds, another natural direction of work studies the computational complexity of finding the best pebbling strategy in RBP. This includes NP-hardness proofs for different variants of the model (RBpebbling2; RBpebbling3; RBpebbling8) and inapproximability results (mpp; RBpebbling3), as well as approximation algorithms for special cases (RBpebbling5; RBpebbling7).
@@
There are also several papers that discuss the idea of generalizing the RBP model to multiple processors or to deeper memory hierarchies (mpp; RBpebbling7; savage1995extending; savage2008unified; elango2014characterizing). However, these generalized settings are typically very complex, and thus the theoretical results in them are limited to concrete DAGs or special cases.
@@
A recent preprint (partial) raises the idea of extending RBP with partial computations. Our work is inspired by this approach, but rather different both in terms of the model and the direction of the results. In particular, the model of (partial) assumes that all nodes have an initialization value that always needs to be loaded from slow memory, motivated by e.g. evaluating polynomials. This incurs an extra I/O step for each sink non-sink node, and hence results in a significantly different (often much higher) I/O cost for most DAGs than either the original RBP or our model. Due to this, most claims in our paper do not carry over to the setting of (partial). In contrast to this, our PRBP model is a direct extension of RBP, where any RBP pebbling strategy translates to a partial-computing PRBP solution of the same I/O cost.
@@
In terms of results, (partial) mainly focuses on DAGs that are not even computable in RBP due to $r\!<\!(\Delta_{in}\!+\!1)$, and shows NP-hardness and approximation methods for the edge case when the cache size is only r=2r\!=\!2. In contrast, we focus on understanding the properties of PRBP in general, and its relation to RBP, for example, how the optimal I/O costs differ between the two models, and which general results from RBP carry over to PRBP.
@@
Regarding the actual model definition, we also change the terminology from (partial) slightly to align it with the original RBP terminology. Specifically, we use “light red” and “dark red” pebbles to denote the two different states of values kept in cache, to signify that these all correspond to red pebbles in the original RBP. Furthermore, the incoming edges of a node that are already aggregated into the final value are not deleted, but rather marked in our case; this indicates that the structure of the computation remains unchanged, and also allows to extend the setting to re-computations.
@@ 3. Pebbling With Partial Computations
The partial-computing red-blue pebble game (PRBP) is defined as follows. Similarly to the original RBP, we have red and blue pebbles, which indicate that the output value of a node $v$ is stored in fast memory and slow memory, respectively. However, in PRBP, we now have two different kinds of red pebbles: a light red pebble on node $v$ indicates that the current value of $v$ is also up-to-date in slow memory, whereas a dark red pebble indicates that the newest value of $v$ is only present in fast memory, and thus, if ever needed in the future, it has to be first saved to slow memory before deletion. The total number of light and dark red pebbles on the DAG is again bounded at any point by the fast memory size $r$. The role of blue pebbles remains unchanged: they still represent values in slow memory, and their number is unlimited.
@@
    partial compute: consider an unmarked edge $(u,v)$ such that: (i) all incoming edges of $u$ are marked, (ii) there is a (light or dark) red pebble on $u$, and (iii) there is either a (light or dark) red pebble on $v$, or no pebble on $v$ at all. Then replace all pebbles on $v$ by a dark red pebble, and mark the edge $(u,v)$.
@@
Similarly to RBP, the initial state only has blue pebbles on the source nodes, and all edges are unmarked. In the terminal state, we now not only require that every sink has a blue pebble, but also that every edge of the DAG is marked; this indicates that all the computations have indeed been fully executed. At any point, the number of red pebbles on the DAG can be at most $r$. As in RBP, the cost of a pebbling is defined as the total number of save and load operations.
Note that in contrast to RBP, this game allows a valid pebbling for any DAG with as few as r=2r\!=\!2 red pebbles, by considering a topological order of $G$ and marking each incoming edge of the next node, with the inputs loaded and outputs saved when necessary.
In our paper, we consider the one-shot version of RBP and PRBP, which means that during a pebbling, the compute rule can be applied only once for every node $v\!\in\!V$ in RBP, and only once for every edge $(u,v)\!\in\!E$ in PRBP. One-shot RBP is a prominent version in the literature (RBpebbling3; RBpebbling7; elango2014characterizing; elango2015characterizing); we focus on this variant because allowing so-called re-computation steps in PRBP would raise further modeling questions. In Section 8.1, we briefly discuss other variants of RBP, and how the PRBP model extends to these settings.
For a given DAG $G$ and choice of $r$, we will use $\texttt{OPT}_{RBP}$ and $\texttt{OPT}_{PRBP}$ to denote the cost of the optimal pebbling strategy in RBP and PRBP, respectively. Regarding notation, we also commonly use $[d]$ to denote the set of integers $\{1,...,d\}$.
@@ 4. Fundamental properties of PRBP
One fundamental question about this extended model is how the partial compute steps affect the optimal I/O cost of a specific computation. Note that we only compare $\texttt{OPT}_{RBP}$ and $\texttt{OPT}_{PRBP}$ for cases when $(\Delta_{in}+1)\leq r$, i.e. a valid pebbling exists in both RBP and PRBP.
@@
Furthermore, note that both in RBP and PRBP, any valid pebbling strategy needs to load every source node at least once, and needs to save every sink node at least once. If the total number of source and sink nodes in $G$ is $t$, then we refer to this value $t$ as the trivial cost, and we have $\texttt{OPT}_{RBP}\geq t$ and $\texttt{OPT}_{PRBP}\geq t$.
@@ 4.1. Motivating examples
Towards a more practical example, consider a fundamental operation in linear algebra, namely, the multiplication $A\!\cdot\!x=y$ of an $m\times m$ matrix $A$ with an $m\times 1$ vector $x$. The corresponding computational DAG consists of $m^{2}\!+m$ source nodes, $m^{2}$ intermediate nodes with in-degree $2$, and $m$ sink nodes of in-degree $m$.
@@
In PRBP, we can devise a strategy that keeps the $m$ partially computed values of the output $y$ in fast memory, and only uses $3$ further red pebbles besides this. In particular, we can iterate through the entries $x_{i}$ of the input vector, first loading $x_{i}$ into fast memory, and then for each row of the matrix, we load the corresponding $A_{j,i}$, multiply it with $x_{i}$, and add this to $y_{j}$. Note that $A_{j,i}$ can be deleted after computing $A_{j,i}\!\cdot\!x_{i}$, and then $A_{j,i}\!\cdot\!x_{i}$ can be deleted after updating $y_{j}$, so this only requires two red pebbles besides the one on $x_{i}$. Afterwards, $x_{i}$ can also be deleted. This strategy comes with only the trivial I/O cost of $\texttt{OPT}_{PRBP}=m^{2}+2m$.
In RBP, we can show that at least one non-trivial I/O step always happens between the computation of any two consecutively computed sink nodes $y_{j_{1}}$ and $y_{j_{2}}$. Note that computing $y_{j_{1}}$ requires $(m+1)$ red pebbles simultaneously, and thus we can have at most (m1)(m-1) red pebbles in the remaining nodes. As such, there is an index $i\in[m]$ such that neither the source node $x_{i}$ nor the intermediate node $A_{i,j_{2}}$ has a red pebble. If $A_{i,j_{2}}$ already has a blue pebble, then this means we have to load it before computing $y_{j_{2}}$. Otherwise, if $A_{i,j_{2}}$ has no pebble, then we need to load $x_{i}$ to compute it; note that $x_{i}$ has been loaded before to compute $A_{i,j_{1}}$, so this is also a non-trivial I/O. Altogether, this gives $\texttt{OPT}_{RBP}\geq m^{2}\!+\!3m-1$. ∎